An IV is infusing at 25 gtt/min with a microdrop infusion set. How many mililiters will have infused over 4.3 hours? Round answer to a whole number. 7. mL Answer
Answer #1

Given values are

Rate = 25 gtt/min

Microdrop means 60 drops in min

Time = 4.3 hours

i.e, =4.3* 60 = 258 minutes

So the equation is,

Rate per min = (total volume to be infused * drop factor)/ time in minutes

That is , 25 =( Volume * 60)/ 258

Volume to be infused =( 25*258)/60

=107.5 ml

Answer is 107.5 ml